OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the second equation is substituted into the first equation. 2x - 7y = 4 3x + y = -17

mathslover (mathslover):

First of all, transpose 3x to the Right Hand Side. \[\large{3x+y=-17}\] \[\large{y = -17-3x}\]

mathslover (mathslover):

Now put the above equation into the first one on the place of y

OpenStudy (anonymous):

All you do is rearrange the second equation to get y by itself. Then, you can plug it in where y is in the first equation. Just subtract 3x over to the other side to get y=-17-3x.

mathslover (mathslover):

@alexislovier what do you get now?

OpenStudy (anonymous):

uhm. i dont know

OpenStudy (anonymous):

Im horrible at math .

mathslover (mathslover):

As the first equation we have : 2x - 7y = 4 and from the above : y = -17-3x Now put this value of y in the first equation : 2x - 7y = 4 \(2x-7y=4\) But y = -17-3x So : \(2x- 7(-17-3x) =4\) Solve it now.

OpenStudy (anonymous):

so you substract 2x and 3x?

OpenStudy (anonymous):

hello?

mathslover (mathslover):

No! : first multiply 7 to the term : (-17-3x)

OpenStudy (anonymous):

From 2x-7(-17-3x)=4, you distribute the -7. So multiply both of the things in the parens by -7. That gives you 2x+119+21x=4. Then, simplify it to 23x+119=4. Subtract 119 from both sides to get 23x=-115. Divide both sides by 23. X is -5.

mathslover (mathslover):

we have : 2x - 7(-17-3x) For solving this : we will do multiply (7) to (-17-3x) .. What do you get after multiplying @alexislovier

OpenStudy (anonymous):

Now, plug -5 back in for x in one of the original equations. Lets use (2x-7y=4). So (2(-5)-7y=4) Simplify and solve for y. That's -10-7y=4. Add -10 to both sides for -7y=14. Now divide both sides by -7 to get y=-2. Your answers are x=-5,y=-2.

OpenStudy (anonymous):

A. 2(3x - 17) - 7y = 4 B. 2(-3x - 17) - 7y = 4 C. 2x - 7(3x - 17) = 4 D. 2x - 7(-3x - 17) = 4

OpenStudy (anonymous):

i have too get one of those

OpenStudy (anonymous):

Oh then you stop at this step. 2x-7(-17-3x)=4.

mathslover (mathslover):

That you have : \(2x - 7(-17-3x) = 4\) which is the last option.

OpenStudy (anonymous):

ohhhhhh ok! got cha

mathslover (mathslover):

Good! also WELCOME TO OPENSTUDY

OpenStudy (anonymous):

can you guys help me with another one ?